OpenStudy (anonymous):

Find the heights of a square pyramid that has a volume of 12 ft.³ and a baseline of 3 feet

OpenStudy (welshfella):

Volume of a pyramid = (1/3) * area of the base * height

OpenStudy (welshfella):

so here we have 12 = (1/3) * 3^2 * h solve for h

OpenStudy (welshfella):

note that the area = 3^2 because the base is a square

OpenStudy (welshfella):

can you solve this?

OpenStudy (koikkara):

\(12 = (1/3) * 3^2 * h\) @katelyn-andrea

OpenStudy (anonymous):

t would be 12= (1/3)•9•h.. But after that step what do i do? @welshfella

OpenStudy (koikkara):

\(12 = (1/3)9 h\) \(12=3h\) \(h=4\)